f(x) = 2x + 4
and
g(x) = 3x - 5
What would (f - g)(2) = ?

Answer by jim_thompson5910(35256)

Since (f - g)(2) = f(2) - g(2), we need the individual values of f(2) and g(2):

f(x) = 2x + 4
f(2) = 2(2) + 4 ... plug in x = 2 and evaluate
f(2) = 8

--------------

g(x) = 3x - 5
g(2) = 3(2) - 5 ... plug in x = 2 and evaluate
g(2) = 1

-------------

So,

(f - g)(2) = f(2) - g(2)
(f - g)(2) = 8 - 1
(f - g)(2) = 7